# M3-20M: A Large-Scale Multi-Modal Molecule Dataset for AI-driven Drug Design and Discovery

### v1

@@ 5.2 Molecular Property Prediction
while $n$ is the number of samples, $y_{i}$ is predicted value, and $\hat{y}_{i}$ is the ground truth value.
The results in Table 6 show that providing more information with multi-modal data in the prompt can help LLMs better understand molecules and their chemical properties. Except for the task of predicting the dipole moment using GPT-4, all other models and tasks achieve the best results on the multi-modal dataset, which proves the effectiveness and high quality of our dataset.
@@
Table 6: Performance comparison of models using different modal data.
Classification tasks:
We evaluate the performance of single-modal and multi-modal datasets on various molecular property prediction tasks, focusing on accuracy mean (ACC Mean), accuracy variance (ACC Variance), and accuracy standard deviation (ACC Standard Deviation). The results are summarized in Table 7.
ACC Mean measures the average accuracy of the models. Multi-modal datasets show improvements over single-modal datasets. For example, the BACE-MM dataset’s ACC Mean increases from 0.5680 to 0.5780, and the BBBP-MM dataset sees an increase from 0.2280 to 0.2720. The ClinTox-MM dataset improves from 0.9260 to 0.9280, while the HIV-MM dataset shows a higher ACC Mean for single-modal data (0.9740) compared to multi-modal (0.9680). These results suggest that multi-modal datasets typically enhance model accuracy. ACC Variance indicates the variability in accuracy. The ClinTox-MM dataset’s variance decreases from 0.0005 to 0.0002 with multi-modal data, indicating more stable performance. ACC Standard Deviation provides insight into the distribution of accuracy values. The ClinTox-MM dataset sees a reduction from 0.0230 to 0.0130, suggesting enhanced consistency.
@@
Table 7: Accuracy comparison between single-modal and multi-modal datasets for molecular property prediction tasks.
We compare the performance of single-modal and multi-modal datasets on various molecular property prediction tasks with the Tox21-MM dataset as the benchmark. The results are summarized in Table 8.
For ACC Mean, multi-modal datasets generally show an improvement. For example, in the NR-AR task, the ACC Mean increases from 0.9620 to 0.9680 with multi-modal data. Similarly, the NR-AhR task shows a rise from 0.8740 to 0.9020, and the NR-Aromatase task improves from 0.9620 to 0.9720. These results suggest that multi-modal datasets generally enhance accuracy. For ACC Variance, both the NR-AR and NR-PPAR-gamma tasks show equal variances of 0.0003 and 0.0002, indicating consistent performance. The NR-ER task significantly reduces variance from 0.0014 to 0.0003 with multi-modal data, highlighting improved stability. For ACC Standard Deviation, the SR-HSE task decreases significantly from 0.0295 to 0.0114 with multi-modal data, suggesting enhanced consistency.
@@
Table 8: Accuracy comparison between single-modal and multi-modal datasets on Tox21-MM.
Additionally, fine-tuning Llama3-8b with our data shows notable improvements. Results in Table 9 shows that fine-tuning Llama3-8b with our M3-20M dataset leads to notable improvements in model performance across various tasks. Specifically, the multi-modal results show significant gains over the single-modal baseline for the BACE, Tox21, HIV, and ClinTox datasets, with the most pronounced improvement seen on the BACE dataset, where performance increases from 0.844 to 0.887. This proves that our dataset has a wide range of application scenarios and practical value. In fact, our dataset can also be used for downstream tasks such as name prediction, reaction prediction, retrosynthesis, text-based molecule design, molecule captioning etc.
Table 9: Fine-tuning Llama3-8b with our M3-20M dataset
@@
BBBP 0.959 0.959
BACE 0.844 0.887
Tox21 0.817 0.849
HIV 0.780 0.803
ClinTox 0.985 0.998
@@ 6 Conclusion
In this paper, we introduce M3-20M, a new and large-scale multi-modal molecular dataset containing over 20 million molecules. Extensive experiments demonstrate M3-20M’s effectiveness in boosting the performance of various models in molecule generation and property prediction. M3-20M’s integration of SMILES strings, 2D graphs, 3D structural data, physicochemical properties, and molecular textual descriptions provides a rich foundation for developing high-performance generative models for AI-driven drug design and discovery.
{acknowledgement}
This work was supported in part by the National Natural Science Foundation of China (NO. 62372326, No. 62172300).
@@ 7 Data and Software Availability Statement
The dataset is openly accessible for non-commercial use, hosted on a publicly available platform (https://github.com/bz99bz/M-3). This project also provides detailed dataset documentation and the workflow code discussed in the present paper.

### v2

@@ 4.3.1 Regression Tasks
where $n$ is the number of samples, $y_{i}$ is predicted value, and $\hat{y}_{i}$ is the ground truth value.
Experimental results. The results in Table 6 show that providing more information with multi-modal data in the prompt can help LLMs better understand molecules and their chemical properties. Except for the task of predicting the dipole moment using GPT-4, all other models and tasks achieve the best results on the multi-modal dataset, which proves the effectiveness and high quality of our dataset.
Table 6: Comparison of molecular property prediction performance using multi-modal and single-modal data (regression task).
@@
4.3.2 Classification Tasks
Experimental details. First, we apply FP-ICL to regression tasks on BACE-MM, BBBP-MM, ClinTox-MM, HIV-MM, and Tox21-MM datasets with single-modal and multi-modal data respectively. Four randomly-selected molecules from each of the above datasets are used, with SMILES strings only for single-modal molecules, and SMILES strings, 3D coordinates and textual descriptions for multi-modal molecules.
Then, we employ parameter-efficient fine-tuning with LoRA on Llama3-8b using our dataset M3-20M. Experiments are run on an RTX4090 with 24GB memory. LoRA hyperparameters include a rank of 4 and alpha of 16. Training lasts a maximum of 200 epochs with early stopping after 15-30 epochs without improvement.
Performance metrics. We perform each experiment five times and report the averaged result and variance, using three metrics including accuracy mean (ACC Mean), accuracy variance (ACC Variance), and accuracy standard deviation (ACC Standard Deviation). These metrics provide comprehensive insight into prediction reliability and consistency across the different molecule datasets.
Experimental results. Results from FP-ICL on classification tasks (Table 7) show using multi-modal data generally outperforms using single-modal data in molecular property prediction. On the majority of datasets including BACE-MM, BBBP-MM and ClinTox-MM, mean accuracy is improved, only on HIV-MM there is a slight performance decrease. Using multi-modal data usually yields reduced variance and standard deviation, particularly on the ClinTox-MM dataset, indicating enhanced prediction stability and consistency. These findings suggest that integrating multiple modalities provides complementary information and leads to more robust and accurate molecular property predictions.
Table 7: Comparison of molecular property prediction performance using multi-modal and single-modal data (classification task).
@@ 4.3.2 Classification Tasks
Results on the Tox21-MM dataset are summarized in Table 8, from which we can still see that using multi-modal data can generally enhance molecular property prediction performance across multiple toxicity subtasks. On 8 of 12 subtasks we witness improved or equivalent accuracy, with notable improvements in NR-AhR and NR-Aromatase tasks. Using multi-modal data also frequently produces more consistent predictions, evidenced by the reduced variance and standard deviation in several subtasks, particularly NR-ER and SR-HSE. These results confirm that complementary information from multiple modalities enhances accuracy and reliability of toxicity prediction.
Table 8: Comparison of molecular property prediction performance using multi-modal and single-modal data on Tox21-MM (classification task).
@@
Table 9 presents the results of fine-tuning Llama3-8b with our M3-20M dataset. We also see notable improvements on model performance across various tasks. Concretely, using multi-modal data achieves significant gain over using single-modal data on BACE-MM, Tox21-MM, HIV-MM, and ClinTox-MM datasets, with the most pronounced improvement on the BACE dataset, where performance increases from 0.844 to 0.887. This shows that our dataset’s value and potential. Our dataset can also be used for other downstream tasks such as name prediction, reaction prediction, retrosynthesis, text-based molecule design, and molecule captioning, etc.
Table 9: Fine-tuning Llama3-8b with our M3-20M dataset demonstrates significant performance improvement in terns of ACC mean.
@@
BBBP-MM 0.959 0.959
BACE-MM 0.844 0.887
Tox21-MM 0.817 0.849
HIV-MM 0.780 0.803
ClinTox-MM 0.985 0.998
4.4 Statistic Analyses and Results
This section presents statistical information on the M3-20M dataset to highlight its comprehensive annotations, diverse textual descriptions, and physicochemical properties, which can significantly enhance its utility for molecular research and AI-driven drug discovery.
4.4.1 Textual Descriptions
The PubChem description bar is relatively small, reaching only 360,133, while the M3-20M annotation description bar towers over it, marking a count of 20,249,090. This enhancement is critical for AI-driven drug design and discovery, as the increased volume and details of descriptions facilitate more comprehensive analysis and model training.
4.4.2 Word Distribution in Description Texts
Figure 5 shows the 20 most frequently occurring words in textual descriptions of our dataset, including “natural”, “product”, “acid”, and “metabolite” etc. These terms appear tens of thousands of times, reflecting the dataset’s emphasis on biologically derived compounds. The frequency distribution of descriptors like “conjugate”, “streptomyces” and “aza macrocycle” reveals the dataset’s focus on microbial secondary metabolites and their derivatives. This lexical analysis provides insight into the composition of the textual descriptions and highlights the predominance of natural product chemistry within the represented molecular space.
Figure 5: 20 most-frequent words in textual descriptions.
Physicochemical Properties Number Reference
Molecular Weight 19174000 Computed by PubChem 2.2
XLogP3 16340203 Computed by XLogP3 3.0
Hydrogen Bond Donor Count 19173687 Computed by Cactvs 3.4.8.18
Hydrogen Bond Acceptor Count 19173559 Computed by Cactvs 3.4.8.18
Rotatable Bond Count 19173464 Computed by Cactvs 3.4.8.18
Exact Mass 19173398 Computed by PubChem 2.2
Monoisotopic Mass 19173327 Computed by PubChem 2.2
Topological Polar Surface Area 19173287 Computed by Cactvs 3.4.8.18
Heavy Atom Count 19173259 Computed by PubChem
Formal Charge 19173222 Computed by PubChem
Complexity 19173210 Computed by Cactvs 3.4.8.18
Isotope Atom Count 19173193 Computed by PubChem
Defined Atom Stereocenter Count 19173170 Computed by PubChem
Undefined Atom Stereocenter Count 19173153 Computed by PubChem
Defined Bond Stereocenter Count 19173147 Computed by PubChem
Undefined Bond Stereocenter Count 19173137 Computed by PubChem
Covalently-Bonded Unit Count 19173132 Computed by PubChem
Compound Is Canonicalized 19173130 Computed by PubChem
Physical Description 2031 CAMEO Chemicals database
Color/Form 246 Hazardous Substances Data Bank (HSDB)
Odor 100 Hazardous Substances Data Bank (HSDB)
Boiling Point 130 CAMEO Chemicals database
Melting Point 633 CAMEO Chemicals database
Flash Point 35 CAMEO Chemicals database
Solubility 10813 CAMEO Chemicals database
Density 170 CAMEO Chemicals database
Figure 6: Visualization of molecular space with uniform manifold approximation and projection (UMAP).
Figure 6 illustrates the diversity of our dataset in a two-dimensional space using UMAP mcinnes2018umap . UMAP (Uniform Manifold Approximation and Projection) excels as a dimensionality reduction technique by preserving both local and global structure while offering computational efficiency. It is particularly valuable for visualizing high-dimensional molecular fingerprints where topological relationships between compounds must be maintained. The UMAP projection of the M3-20M dataset reveals a predominantly interconnected chemical space with a dense central region and several peripheral clusters, suggesting substantial molecular diversity within a coherent chemical landscape.
4.4.3 Dataset Maintenance
A six-member maintenance team (three doctoral and three master’s students) will perform bi-weekly updates to the M3-20M dataset, correcting errors and implementing updates. This protocol enhances the dataset’s reliability for molecular representation research and supports more robust model training.
@@ 5 Conclusion
In this paper, we introduce M3-20M, a new and large-scale multi-modal molecule dataset containing over 20 million molecules. Extensive experiments demonstrate M3-20M’s effectiveness in boosting the performance of various models in molecule generation and property prediction. M3-20M’s integration of SMILES strings, 2D graphs, 3D structural data, physicochemical properties, and molecular textual descriptions provides a rich source for developing high-performance generative models for AI-driven drug design and discovery.
@@ 6 Data and Software Availability Statement
The dataset is openly accessible under the GNU General Public License v3.0, hosted on a publicly available platform (https://github.com/bz99bz/M-3).
7 Author contributions statement
Siyuan Guo and Shuigeng Zhou conceived the experiment(s), Siyuan Guo, Lexuan Wang, Chang Jin, Jinxian Wang, Han Peng, and Huayang Shi conducted the experiment(s), Siyuan Guo, Lexuan Wang, and Chang Jin analysed the results. Siyuan Guo, Wengen Li, Jihong Guan, and Shuigeng Zhou wrote and reviewed the manuscript.
8 Acknowledgments
The authors thank the anonymous reviewers for their valuable suggestions. This work was supported in part by the National Natural Science Foundation of China (NO. 62372326, No. 62172300).
Key Points M3-20M is the largest open-access multi-modal molecule dataset for AI-driven drug design and discovery, with over 20 million molecules. M3-20M has comprehensive modalities, integrating SMILES, 2D/3D structures, physicochemical properties, and textual descriptions. Experiments show that M3-20M can significantly boost the model performance in molecular generation and property prediction tasks. M3-20M supports diverse downstream tasks, including molecule generation, molecular property prediction, lead optimization, and virtual screening, etc.